# Insights into the Incorporation of Signal Information in Binaural Signal Matching with Wearable Microphone Arrays

### v2

@@ Abstract
The increasing popularity of spatial audio in applications such as teleconferencing, entertainment, and virtual reality has led to the recent developments of binaural reproduction methods. However, only a few of these methods are well-suited for wearable and mobile arrays, which typically consist of a small number of microphones. One such method is binaural signal matching (BSM), which has been shown to produce high-quality binaural signals for wearable arrays. However, BSM may be suboptimal in cases of high direct-to-reverberant ratio (DRR) as it is based on the diffuse sound field assumption. To overcome this limitation, previous studies incorporated sound-field models other than diffuse. However, this approach was not studied comprehensively. This paper extensively investigates two BSM-based methods designed for high DRR scenarios. The methods incorporate a sound field model composed of direct and reverberant components. The methods are investigated both mathematically and using simulations, finally validated by a listening test. The results show that the proposed methods can significantly improve the performance of BSM , in particular in the direction of the source, while presenting only a negligible degradation in other directions. Furthermore, when source direction estimation is inaccurate, performance of these methods degrade to equal that of the BSM, presenting a desired robustness quality.
@@ I Introduction
In an effort to overcome the limitations of existing beamforming-based methods and to accurately reproduce binaural signals captured by arrays of arbitrary geometry, a number of techniques have been developed [9, 10, 11, 12]. The binaural signal-matching (BSM) technique is one such technique [13, 43]. BSM entails estimating the binaural signals directly from the array measurements using least-squares (LS) optimization. To enhance perception, at high frequencies, Magnitude Least-Squares (MagLS) optimization is used instead [26]. A BSM design was recently presented for a semi-circular array with varying numbers of microphones, incorporating head tracking [13]. The study revealed that BSM accuracy heavily depends on microphone positioning, leading to poor performance when a listener’s ear is distant from the recording microphones. In addition, the BSM which is designed under a diffused sound-field assumption may be sub-optimal under high Direct-to-Reverberant Ratio (DRR) conditions, for example [29].
@@
Parametric spatial audio was also investigated as an alternative for binaural reproduction. In these approaches, the sound field is decomposed into direct and reverberant components, each estimated and reproduced independently [21, 22, 23, 24, 25]. While this approach has shown promising results, its accuracy depends on the estimation of signal-dependent model parameters, such as the direction-of-arrival (DOA) of sound sources, and the direct-to-reverberant ratio (DRR) of the sound field. This drawback of the parametric approach, and the limitations of the beamforming approach detailed above, define the current gaps in binaural reproduction using arrays of arbitrary configurations, highlighting the need for improved solutions.
@@
This paper aims to address the shortcomings of current methods by exploring the incorporation of signal information into BSM. The potential benefit would be high stability and fidelity due to the nature of the BSM linear filtering, with improvement in performance at high frequency and due to head rotation delivered by modeling sound sources. Two relevant approaches have been recently proposed. The first is to incorporate signal information in the correlation matrix employed in the BSM design [29], and the second is based on parametric spatial coding, e.g. COMPASS [25], where the residual sound field complementing individual source modeling is reproduced using a BSM like approach. While both approaches showed promising potential, it is unclear from current literature which is preferred, or in general, what is the best way to incorporate signal information into BSM.
This paper presents a comprehensive theoretical formulation and experimental investigation to cover this gap. The work is a significant extension of a conference paper by the same authors that investigated the feasibility of incorporating signal information into BSM [44]. The work shows both methods exhibit commendable performance when the array microphones are in close proximity to the ears of a listener virtually positioned in the sound scene. However, the performance degrades as the microphones move further from the listener ears. Furthermore, a novel performance metric is employed to analyze the expected relative errors associated with each approach across different directions. The main contributions of this paper are:
@@
    A new performance metric of BSM is proposed. This metric quantifies the expected relative error as a function of sound source direction. The metric is referred to as the “directional error”. Utilizing this metric, it becomes evident that the BSM exhibits distinct errors that are dependent on the DOA of the sound sources, highly related to HRTF magnitude. This behavior is further validated using interaural time difference (ITD) and interaural level difference (ILD) analysis.
@@
    The presentation, analysis, and comparison of two distinct approaches that incorporate signal information into the BSM method, in particular information on source DOA. These methods are shown to overcome the limitations imposed by the directional behavior of the BSM, performing well for all source directions. This behavior is also validated through ITD and ILD analysis.
@@
    Furthermore, performance analysis of the methods reveals a discernible trade-off, wherein the enhancement of BSM performance in the DOA of a sound source is counterbalanced by only a marginal decrement in performance across most other directions. In particular, one method performs better in directions away from the source. This method was also found to be more robust to errors in the estimation of the source DOA.
  4. 4.
    Finally, a comprehensive simulation study and a listening experiment are performed that substantiate the accuracy and quality of the different methods with a wearable microphone array, validating the methods behavior and limitations.
@@ II-B Binaural Signal Matching
Minimizing the error in Eq. (6) leads to the following solution:
$\mathbf{c}^{l,r}_{opt}(f)=(\mathbf{V}\mathbf{R_{s}}\mathbf{V}^{H}+\mathbf{R_{n}})^{-1}\mathbf{V}\mathbf{R_{s}}[\mathbf{h}^{l,r}]^{*}$ (7)
@@
$\mathbf{c}^{l,r}_{BSM}(f)=(\mathbf{V}\mathbf{V}^{H}+\frac{1}{SNR}\mathbf{I}_{M})^{-1}\mathbf{V}[\mathbf{h}^{l,r}]^{*}$ (8)
@@ III BSM Approaches that Incorporate Signal Information
$\mathbf{x}(f)=\mathbf{V_{d}}\mathbf{s_{d}}+\mathbf{V_{r}}\mathbf{s_{r}}+\mathbf{n}$ (9)
@@ IV-A Normalised Mean-Square Error
IV-B Directional Error
This subsection introduces a novel measure to analyze the BSM method, referred to as the directional error. Additionally, a methodology for computing this error is provided. For simplicity, throughout this section the dependency on the frequency is omitted. Assume the BSM filter $\mathbf{c}$ is computed as outlined in Section II-B, and that the sound-field described in Section II-A consists solely of a single source, having a DOA denoted as $\Omega$. Additionally, assume that the source signal $s$ is wide sense stationary. Next, the directional error at frequencies below 1.51.51.5\,1.5kHz is defined as the normalized mean square error, as a function of the source signal DOA. This error can be calculated under these assumptions using Eq. (4), further assuming no noise and finally applying normalization, as follows:
$err^{l,r}_{dir}(\Omega)=\begin{cases}\frac{\|\mathbf{v}(\Omega)^{H}\mathbf{c}^{l,r}-[{h}^{l,r}(\Omega)]^{*}\|_{2}}{|{h}^{l,r}(\Omega)|},&\text{for }f<1.5\text{kHz}\\ \frac{\||[\mathbf{v}(\Omega)]^{H}\mathbf{c}^{l,r}|-|{h}^{l,r}(\Omega)|\|_{2}^{2}}{|{h}^{l,r}(\Omega)|},&\text{for }f\geq 1.5\text{kHz}\end{cases}$ (26)
where $err^{l,r}_{dir}(\Omega)$ refers to the directional error for the left and right channels, $\mathbf{c}^{l,r}$ is the BSM filter, ${h}^{l,r}(\Omega)$ refers to the HRTF with direction $\Omega$, and $\mathbf{v}(\Omega)$ is a $M\times 1$ complex array steering vector corresponding to a given DOA. For high-frequency signals, due to consideration of sound perception [39], the directional error is based on signal magnitude.
IV-C ITD and ILD
@@ V Simulation Study
In this section, a comparison between the two approaches, presented in Section III, and the standard BSM, presented in Section II-B, is provided. Broadly, this section aims to study the benefits and to gain insights into signal-dependent BSM approaches, in particular studying direction-dependent performance highlighting the points made in Section III-C regarding the methods modeling detail and robustness to error.
@@ V-A Simulation Set-Up
In this section, a semi-circular microphone array, mounted on a rigid sphere, comprising $M=6$ omnidirectional microphones arranged on the horizontal plane was simulated. The semi-circular array was chosen since it is a simplified representation of a head-worn array, such as augmented reality glasses. The microphone positions were denoted using spherical coordinates $(r,\frac{\pi}{2},\phi_{m})$ for $m=1,\ldots,M$ relative to the array center, where $r=10\,$cm and $\phi_{m}=\frac{\pi}{2}-\frac{\pi(m-1)}{M-1}\,$rad. The steering vectors of this array were computed analytically in the SH domain following section 4.2 in [16] using the array geometry and the DOAs of the far-field sources modeled as plane waves. The HRTF in the simulation was obtained from the Cologne database measurements of the Neumann KU100 manikin [17]. The HRTFs for the DOAs of the assumed sources were interpolated in the SH domain using a SH order of 30. The assumed listener head was aligned with the positive $x$ axis. An illustration of the array position relative to the head position and alignment is presented in Figure 1. An orientation of $90^{\circ}$ as shown in Figure 1 of the array relative to the head was found to be the most challenging for the BSM algorithm [20] since the right ear is relatively far from any of the arrays microphones compared to the standard case of no head rotation as shown in 1.
@@ V-B Methodology
Following the computation of the three filters, the binaural signals for all filters were computed using Eq. (3) by replacing the appropriate filter $\mathbf{c}^{l,r}$. The reference binaural signals, were calculated by convolving the HRTFs of the left and right ears with the HOA signals of order 14 that were calculated using the image method as described in section V-A. Finally, the errors measures as presented in Section IV were computed using the binaural signals for the three methods, and the reference binaural signals.
V-C BSM Directional-Error
This section presents the NMSE of the standard BSM method, as defined in Section IV-A, assuming a diffuse sound field as detailed in [20]. Figure 2 presents this NMSE as a function of frequency for three different head rotations as described in Section V-B.
As can be observed in Figure 2, the error is characterised by a lower errors in low frequencies compared to high frequencies and by a higher error for that ear that is further from the array, i.e. the right ear. This error presents the overall error due to a diffuse sound field and does not provide information on the directional characteristics of the BSM method.
Figure 3: Surface plots illustrating the directional error of the BSM method, with the left ear presented on the left and the right ear on the right. The directional error is computed in accordance with Section IV-B, considering a consistent elevation of $90^{\circ}$. The frequency is represented on the y-axis, while the azimuth angle $\phi$ in degrees is depicted on the x-axis, assuming no head rotation.
Figure 4: Surface plots of the magnitude of the HRTF using a flipped color-map, assuming a constant elevation of $90^{\circ}$, where the y axis is the frequency and the x axis is the azimuth angle $\phi$ in degrees. The HRTF in the simulation was obtained from the Cologne database [17] measurements of the Neumann KU100 manikin. The HRTFs for the DOAs of the assumed sources were interpolated in the SH domain using a SH order of 30.
In order to better understand the directional behavior of the BSM method, the directional error, as defined in Section IV-B, is presented in this section. Figures 3 presents the directional error on the horizontal plane assuming an elevation of $90^{\circ}$ as a function of frequency for the BSM method assuming no head rotation as described in Section V-B. An abrupt change in the frequency of $1.5$ KHz can be observed as this is the MagLS cutoff frequency described in Section V-B. It can also be observed that the error changes quite significantly as a function of direction and frequency. While the error is consistently low below about 1kHz, it changes significantly with azimuth direction at higher frequencies. Figure 4 presents the HRTF magnitude as a function of direction and frequency using a flipped color map compared to Figures 3. It can be observed that the error in Figures 3 resembles the HRTF magnitude in Figure 4 when using this flipped color-map. This means that the error is high at directions where the HRTF magnitude is small. This behavior may become a limitation suggesting higher error at the far ear due to the head shadowing effect.
Figure 5: Surface plots illustrating the directional error of the BSM method, with the left ear presented on the left and the right ear on the right. The directional error is computed in accordance with Section IV-B, considering a consistent elevation of $90^{\circ}$. The frequency is represented on the y-axis, while the azimuth angle $\phi$ in degrees is depicted on the x-axis, assuming a $50^{\circ}$ head rotation.
Figures 5 presents the directional error on the horizontal plane assuming an elevation of $90^{\circ}$ for various frequencies for the BSM method assuming a $50^{\circ}$ head rotation. In the case of $50^{\circ}$ head rotation the right ear moves away from the array while the left ear remains relatively close to the array as illustrated in Figure 1. The directional error of the left ear does not change by much compared to Figures 3, while a clear increase in the directional error can be observed for the right ear, as further illustrated next.
Figure 6: Surface plots illustrating the increase in error per direction of the BSM method with a $50^{\circ}$ head rotation compared to no rotation, i.e. the difference between Figure 5 and Figure 3. The left ear is presented on the left, and the right ear on the right, both maintaining a consistent elevation of $90^{\circ}$. The y-axis corresponds to frequency, while the x-axis represents the azimuth angle $\phi$ in degrees.
Figures 6 presents the difference between Figures 5 and 3. Upon examination of Figures 6, it becomes evident that the increase in NMSE error associated with the head rotation of $50^{\circ}$, as illustrated in Figure 2, exhibits a directional pattern. Additionally, a closer look at Figure 4 and 6 reveals that the predominant portion of the NMSE error stemming from head rotation, can be linked to directions characterized by small HRTF magnitudes.
V-D Analysis of ITD and ILD
@@
Figure 7: The ITD (a) and ILD (b) evaluated using the direct sound component of scenario 1 in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-C. No head rotation was employed.
@@
Performance metrics studied thus far are rooted in the Mean Squared Error (MSE). However, this might not fully capture human perception. Therefore, a study employing perceptually driven measures, ITD and ILD, is introduced. The ILD and ITD, as outlined in Section IV-C, are presented in this section, for scenarios detailed in Table I comparing the different methods.
@@
The calculation of the ILD and ITD can be divided into two parts. First, a source in a room was simulated generating a scenario in Table I. The recorded signals at the array were then used to compute filters for the methods as described in Section V-B. Now, using the same filters, binaural signals were computed but this time for a sound field composed of a single source in free-field, with the same direction of the source in the room in Table I, and ILD and ITD were calculated as defined in Section IV-C, for this specific source direction. Finally, this same process of computing filters and then ILD and ITD, was repeated for a range of source directions to complete the ILD and ITD analysis. This somewhat non-standard analysis was employed so that the filters could be computed using realistic signals, but then evaluated in terms ILD and ITD in the direction of the source only, as the direct sound is known to dominate sound localization.
@@
Figure 8(a) illustrates the performance of ILD and ITD under the same conditions of Figure 7, but with $50^{\circ}$ head rotation. The figure shows that ILD and ITD error increased significantly for the BSM method and insignificantly for the other two methods. Furthermore, it is noticeable that the error did not uniformly increase across all directions, demonstrating the directional NMSE performance elucidated in Section V-C. This shows the superior performance of the d-BSM and COM methods in terms of ITD and ILD when considering head rotation. This can be explained by the explicit information of the source direction embedded in these two methods.
@@
Figure 8: The ITD on the left and ILD on the right evaluated using the direct sound component of the three different scenarios described in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-C. A $50^{\circ}$ head rotation was employed.
@@
V-E Robustness to DOA Estimation Errors
@@
Figure 9: The ITD (a) and ILD (b) evaluated using the direct sound component of scenario 1 in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-C. An error of ${\Omega_{err}}=(\phi=10^{\circ},\theta=0^{\circ})$ and a $50^{\circ}$ head rotation were employed.
@@
V-F Performance at directions away from source DOA
In this subsection, the ILD and ITD performance of the different methods will be calculated for a specific case. In this case, a single source DOA is used for computing the filters, but analysis is performed in all azimuth directions, i.e. also away from source DOA. This analysis is important since it compares the performance of the different methods in directions other than source DOA, which may also be important, e.g. due to an onset of a source in these directions. The calculation of the ILD and ITD in this chapter can also be divided into two parts. First, a source in a room was simulated generating scenario 1 in Table I. The recorded signals at the array were then used to compute filters for the methods as described in Section V-B. In this case, the source was placed at ${\Omega}=(40^{\circ},90^{\circ})$ relative to the array. Now, using the same filters, binaural signals were computed and ILD and ITD were calculated as defined in Section IV-C, for sound sources in free-field scanning all azimuth directions, and an elevation of $90^{\circ}$.
Figure 10: The ITD (a) and ILD (b) of scenario 1 in Table I, with the direct source positioned at a single direction of $40^{\circ}$. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-C. $50^{\circ}$ head rotation was employed.

### v3

# Performance and Robustness of Signal-Dependent vs. Signal-Independent Binaural Signal Matching with Wearable Microphone Arrays

@@ Abstract
The increasing popularity of spatial audio in applications such as teleconferencing, entertainment, and virtual reality has led to the recent developments of binaural reproduction methods. However, only a few of these methods are well-suited for wearable and mobile arrays, which typically consist of a small number of microphones. One such method is binaural signal matching (BSM), which has been shown to produce high-quality binaural signals for wearable arrays. However, BSM may be suboptimal in cases of high direct-to-reverberant ratio (DRR) as it is based on the diffuse sound field assumption. To overcome this limitation, previous studies incorporated sound-field models other than diffuse. However, performance may be sensitive to signal estimation errors. This paper aims to provide a systematic and comprehensive analysis of signal-dependent vs. signal-independent BSM, so that the benefits and limitations of the methods become clearer. Two signal-dependent BSM-based methods designed for high DRR scenarios that incorporate a sound field model composed of direct and reverberant components are investigated mathematically, using simulations, and finally validated by a listening test, and compared to the signal-independent BSM. The results show that signal-dependent BSM can significantly improve performance, in particular in the direction of the source, while presenting only a negligible degradation in other directions. Furthermore, when source direction estimation is inaccurate, performance of of the signal-dependent BSM degrade to equal that of the signal-independent BSM, presenting a desired robustness quality.
@@ I Introduction
In an effort to overcome the limitations of existing beamforming-based methods and to accurately reproduce binaural signals captured by arrays of arbitrary geometry, a number of techniques have been developed [9, 10, 11, 12, 47]. The binaural signal-matching (BSM) technique is one such technique [13, 43]. BSM entails estimating the binaural signals directly from the array measurements using least-squares (LS) optimization. To enhance perception, at high frequencies, Magnitude Least-Squares (MagLS) optimization is used instead [26]. A BSM design was recently presented for a semi-circular array with varying numbers of microphones, incorporating head tracking [13]. The study revealed that BSM accuracy heavily depends on microphone positioning, leading to poor performance when a listener’s ear is distant from the recording microphones. In addition, the BSM which is designed under a diffuse sound-field assumption may be sub-optimal under high Direct-to-Reverberant Ratio (DRR) conditions, for example [29].
@@
Parametric spatial audio was also investigated as an alternative for binaural reproduction. In these approaches, the sound field is decomposed into direct and reverberant components, each estimated and reproduced independently [21, 22, 23, 24, 25]. While this approach has shown promising results, its accuracy depends on the estimation of signal-dependent model parameters, such as the direction-of-arrival (DOA) of sound sources, and the direct-to-reverberant ratio (DRR) of the sound field. This uncertainty in the performance of the parametric approach defines a current gap in BSM-based binaural reproduction, where the choice of what is preferable - the robustness of signal-independent BSM or the performance of signal-dependent BSM — may not be entirely clear.
@@
This paper aims to address this gap by systematically investigating the incorporation of signal information into BSM. To diversify the study, two signal-dependent methods are investigated. The first incorporates signal information in the correlation matrix employed in the BSM design [29], and the second is based on parametric spatial coding, e.g. COMPASS [25], where the residual sound field complementing individual source modeling is reproduced using a BSM like approach.
Performance of these methods is investigated and compared to signal-independent BSM under a wide range of conditions that include a speaker in a reverberant room, with a range of source directions, direction estimations errors, and listener head rotations. The simulation study based on a semi-circular array mounted on a rigid sphere, is finally complemented by a listening test. Important conclusions arise from this study that can be useful when incorporating these methods:
@@
    As expected, a clear improvement in performance is observed for the signal-dependent BSM methods, in particular under head rotations, where one ear is typically rotated away for the array.
@@
    The improved performance at the source direction comes at only a marginal cost of a small reduction in performance across other directions.
@@
    When source direction estimation is poor, the signal-dependent approaches degrade in performance but only to match that of the signal-independent BSM.
These results reinforce the confidence in the signal-dependent BSM approaches, showing relative robustness by not degrading below the signal-independent BSM under conditions of error in parameter estimation.
Note: the work is a significant extension of a conference paper by the same authors that investigated the feasibility of incorporating signal information into BSM [44].
@@ II-B Binaural Signal Matching
Minimizing the error in Eq. (6) leads to the following solution [43]:
$\mathbf{c}^{l,r}_{opt}=(\mathbf{V}\mathbf{R_{s}}\mathbf{V}^{H}+\mathbf{R_{n}})^{-1}\mathbf{V}\mathbf{R_{s}}[\mathbf{h}^{l,r}]^{*}$ (7)
@@
$\mathbf{c}^{l,r}_{BSM}=(\mathbf{V}\mathbf{V}^{H}+\frac{1}{SNR}\mathbf{I}_{M})^{-1}\mathbf{V}[\mathbf{h}^{l,r}]^{*}$ (8)
@@ III BSM Approaches that Incorporate Signal Information
$\mathbf{x}=\mathbf{V_{d}}\mathbf{s_{d}}+\mathbf{V_{r}}\mathbf{s_{r}}+\mathbf{n}$ (9)
@@ IV-A Normalised Mean-Square Error
IV-B ITD and ILD
@@ V Simulation Study
In this section, a comparison between the two approaches, presented in Section III, and the standard BSM, presented in Section II-B, is provided. Broadly, this section aims to study the benefits and to gain insights into signal-dependent BSM approaches under various conditions, including robustness to estimation error.
@@ V-A Simulation Set-Up
In this section, a semi-circular microphone array, mounted on a rigid sphere, comprising $M=6$ omnidirectional microphones arranged on the horizontal plane was simulated. The semi-circular array was chosen since it is a simplified representation of a head-worn array, such as augmented reality glasses [48]. The microphone positions were denoted using spherical coordinates $(r,\frac{\pi}{2},\phi_{m})$ for $m=1,\ldots,M$ relative to the array center, where $r=10\,$cm and $\phi_{m}=\frac{\pi}{2}-\frac{\pi(m-1)}{M-1}\,$rad. The steering vectors of this array were computed analytically in the SH domain following section 4.2 in [16] using the array geometry and the DOAs of the far-field sources modeled as plane waves. The HRTF in the simulation was obtained from the Cologne database measurements of the Neumann KU100 manikin [17]. The HRTFs for the DOAs of the assumed sources were interpolated in the SH domain using a SH order of 30. The assumed listener head was aligned with the positive $x$ axis. An illustration of the array position relative to the head position and alignment is presented in Figure 1. An orientation of $90^{\circ}$ as shown in Figure 1 of the array relative to the head was found to be the most challenging for the BSM algorithm [20] since the right ear is relatively far from any of the arrays microphones compared to the standard case of no head rotation as shown in 1.
@@ V-B Methodology
Following the computation of the three filters, the binaural signals for all filters were computed using Eq. (3) by replacing the appropriate filter $\mathbf{c}^{l,r}$. The reference binaural signals were obtained by convolving the HRTFs of the left and right ears with the HOA signals of order 14, which were computed using the image method as described in Section V-A. Finally, the errors measures as presented in Section IV were computed using the binaural signals for the three methods, and the reference binaural signals.
V-C Analysis of ITD and ILD
@@
Figure 3: The ITD (a) and ILD (b) evaluated using the direct sound component of scenario 1 in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-B. No head rotation was employed.
@@
Performance metrics studied thus far are rooted in the Mean Squared Error (MSE). However, this might not fully capture human perception. Therefore, a study employing perceptually driven measures, ITD and ILD, is introduced. The ILD and ITD, as outlined in Section IV-B, are presented in this section, for scenarios detailed in Table I comparing the different methods.
@@
The calculation of the ILD and ITD can be divided into two parts. First, a source in a room was simulated generating a scenario in Table I. The recorded signals at the array were then used to compute filters for the methods as described in Section V-B. Now, using the same filters, binaural signals were computed but this time for a sound field composed of a single source in free-field, with the same direction of the source in the room in Table I, and ILD and ITD were calculated as defined in Section IV-B, for this specific source direction. Finally, this same process of computing filters and then ILD and ITD, was repeated for a range of source directions to complete the ILD and ITD analysis. This somewhat non-standard analysis was employed so that the filters could be computed using realistic signals, but then evaluated in terms ILD and ITD in the direction of the source only, as the direct sound is known to dominate sound localization.
@@
Figure 4(a) illustrates the performance of ILD and ITD under the same conditions of Figure 3, but with $50^{\circ}$ head rotation. The figure shows that ILD and ITD error increased significantly for the BSM method and insignificantly for the other two methods. This shows the superior performance of the d-BSM and COM methods in terms of ITD and ILD when considering head rotation. This can be explained by the explicit information of the source direction embedded in these two methods.
@@
Figure 4: The ITD on the left and ILD on the right evaluated using the direct sound component of the three different scenarios described in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-B. A $50^{\circ}$ head rotation was employed.
@@
V-D Robustness to DOA Estimation Errors
@@
Figure 5: The ITD (a) and ILD (b) evaluated using the direct sound component of scenario 1 in Table I. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-B. An error of ${\Omega_{err}}=(\phi=10^{\circ},\theta=0^{\circ})$ and a $50^{\circ}$ head rotation were employed.
@@
V-E Performance at directions away from source DOA
In this subsection, the ILD and ITD performance of the different methods will be calculated for a specific case. In this case, a single source DOA is used for computing the filters, but analysis is performed in all azimuth directions, i.e. also away from source DOA. This analysis is important since it compares the performance of the different methods in directions other than source DOA, which may also be important, e.g. due to an onset of a source in these directions. The calculation of the ILD and ITD in this chapter can also be divided into two parts. First, a source in a room was simulated generating scenario 1 in Table I. The recorded signals at the array were then used to compute filters for the methods as described in Section V-B. In this case, the source was placed at ${\Omega}=(40^{\circ},90^{\circ})$ relative to the array. Now, using the same filters, binaural signals were computed and ILD and ITD were calculated as defined in Section IV-B, for sound sources in free-field scanning all azimuth directions, and an elevation of $90^{\circ}$.
Figure 6: The ITD (a) and ILD (b) of scenario 1 in Table I, with the direct source positioned at a single direction of $40^{\circ}$. The orange line represents the BSM method, the purple line corresponds to the d-BSM approach, and the green line depicts the COM approach. ITD and ILD errors are also presented, as defined in Section IV-B. $50^{\circ}$ head rotation was employed.